# Leveraging Large Language Models for Wireless Symbol Detection via In-Context Learning

### v1

@@ Abstract
But whether and how LLMs can benefit challenging non-language tasks in wireless systems is unexplored.
@@ I-A Context and Motivation
At the same time, Large Language Models (LLMs) have been effectively utilized to address a variety of Natural Lanuage Processing (NLP) and Computer Vision (CV) tasks [13, 14]. Despite significant advancements in NLP and CV, pre-trained LLMs have faced limitations in their development within non-linguistic tasks, let alone wireless tasks. Therefore, combining wireless communications and natural language remains a challenge to utilize these capabilities.
In this work, we strive to achieve the best of both worlds by leveraging in-context learning abilities of LLMs on the wireless symbol demodulation task. Our contributions are summarized below:
@@ I-B Related Work
Recently, LLMs such as GPT-3 [14] have showcased the capability of in-context learning. This feature enables a model to generate suitable outputs for a given query input by leveraging a prompt containing input-output example pairs tailored to the task at hand. ICL has proven to be highly effective in linguistic tasks with limited data, as it operates without the need for explicit training. However, ICL performance fluctuates across various prompt templates due to inadequate calibration [15]. In our work, we employ advanced LLM calibration methods, as proposed in [16, 15], to achieve high performance on wireless data while ensuring precise calibration of the LLMs.
@@ II-C Proposed Approach: LLM-based ICL (LMIC)
$P(x_{\rm test},(x_{i},y_{i})_{i=1}^{N})\triangleq d_{1}\oplus d_{2}\oplus\dots\oplus d_{N}\oplus f_{x}(x_{\rm test}),$ (4)

### v2

@@ Abstract
But how LLMs can benefit challenging non-language tasks in wireless systems is not fully unexplored.
@@ I-A Context and Motivation
At the same time, despite significant advancements of Large Language Models (LLMs) in Natural Language Processing (NLP) and Computer Vision (CV) [13, 14], pre-trained LLMs have faced limitations in their development within non-linguistic tasks, let alone wireless tasks. Therefore, combining wireless communications and natural language remains a challenge to utilize these capabilities.
In this work, we aim to achieve the best of both worlds by leveraging in-context learning abilities of LLMs on the symbol detection task. We summarize contributions below:
@@ I-B Related Work
Recently, LLMs such as GPT-3 [14] have showcased the capability of in-context learning. This feature enables a model to generate suitable outputs for a given query input by leveraging a prompt containing input-output example pairs tailored to the task at hand. ICL has proven to be highly effective in linguistic tasks with limited data, as it operates without the need for explicit training. However, ICL performance fluctuates across various prompt templates due to inadequate calibration [15]. Recent studies have shown the potential of using transformer-type sequence models for MIMO detection tasks [23, 24, 25]. However, different from these works [23, 24, 25] that train a sequence model for wireless tasks and then employ ICL, we employ ICL directly on the publically available LLMs in its organic form and use advanced LLM calibration methods, as proposed in [16, 15]. This achieves high performance on the symbol detection problem while ensuring precise calibration of the LLMs.
@@ II-C Proposed Approach: LLM-based ICL (LMIC)
$\!\!\!P(x_{\rm test},(x_{i},y_{i})_{i=1}^{N})\triangleq d_{1}\oplus d_{2}\oplus\dots\oplus d_{N}\oplus f_{x}(x_{\rm test})$ (4)